COLLISIONAL QUENCHING OF OH A^{2}Σ^{+} BY H_{2}: LIFETIMES AND METHODS FOR PROBING THE NONREACTIVE PRODUCT CHANNEL

\underline{LOGAN P. DEMPSEY}, PATRICIA A. CLEARY, CRAIG MURRAY, MARSHA I. LESTER, Department of Chemistry, University of Pennsylvania, Philadelphia, PA, 19104.

Collisional quenching of electronically excited OH A^{2}Σ^{+} radicals by molecular hydrogen is known to be an efficient process. Quenching can proceed through a reactive channel, producing water and atomic hydrogen, or through a nonreactive channel. The goal of this work is to study the outcome of the nonreactive quenching channel, and thereby extract information about the dynamical pathways that lead to quenching. A pump-probe scheme is utilized to determine the OH X^{2}Π population distribution following collisional quenching in a pulsed supersonic expansion. The pump laser excites OH A^{2}Σ^{+} (v′ = 0, N′ = 0), resulting in a significantly reduced lifetime due to quenching. The probe laser monitors the OH X^{2}Π (v′′, J′′) population via laser-induced fluorescence (LIF) on various A − X transitions. To convert observed LIF intensities to a population distribution, it is necessary to account for the fluorescence quantum yield and lifetime of the emitting state. Quenching lifetimes for OH A^{2}Σ^{+} radicals are measured at a temperature of ∼50 K and fit to a single exponential decay. Lifetimes of OH A^{2}Σ^{+} (v′ = 0, N′) show a dramatic increase with rotational level N′, corresponding to a decrease in quenching efficiency. The experimental lifetimes are used to extract pseudo first-order rate constants for quenching as a function of N′ under these conditions.